# Compositional Generalization Requires More Than Disentangled Representations

### v1

@@ 4 Encouraging Compositional Generalization
Figure 3: Inducing compositional generalization through architecture and regularization. (a) schematically illustrates the explicitly compositional architecture with factorized initial embedding layers for regularization (with 1-hot input encodings). (b) show sampled embedding activations corresponding to $x=14$ and $y=14$ for networks trained without and with regularization, respectively. (c) show the volume metric comparison between networks trained without and with regularization as a function of layer depth, respectively. (d) shows the OOD vs. ID MSE plot for various ablation studies over many runs.
Figure 4: Inducing generalizable composition through training on independent factors. (a) show sample grayscale images of 1D Gaussian “stripes” at $x=14$ and $y=14$. (b) shows the generated OOD output as a function of number of Gaussian bumps included in the training dataset. (c) shows the data efficiency scaling of the stripes-only ($2N$) vs. bumps-only ($N^{3}$) to reach $90\%$ accuracy of $x$ and $y$ generation as a function of image size $N$ (accuracy assessed based on the location of the darkest pixel). Here stripes-only dataset has $2N$ scaling due to its ability to compositionally generalize zero-shot. (d) show the volume metric as a function of layer depth of the network trained on a dataset of stripes + bumps. (e) shows the 2D neuron activations across different channels in the first layer (layer 0) of a network trained trained on a dataset of stripes + 50 bumps. (f) shows neural tuning curves of two sample neurons at layer 0 and layer 3 of the same network as in (e).
@@
4.1 Architecture and Regularization For Low-Rank Embeddings
One approach is to architecturally encourage an expansion of the abstract latent representations into 2D embedding filters, while applying explicit rank regularization during the learning of these expansions (illustrated in Fig. 3(a)). For example, Cichocki et al. (2009) show that applying tensor factorization constraints can reduce spurious cross-factor interactions. We add a matrix/tensor factorization penalty to a set of 2D embedding filters for each feature dimension, encouraging them to remain rank-limited. Briefly, to enforce factorization in each embedding matrix $W_{i,j}$, we adopt an SVD-like decomposition to each 2D input embedding filter:
@@ 4.1 Architecture and Regularization For Low-Rank Embeddings
where $r$ is an upper bound on the rank, $\lambda^{(n)}$ are scalar coefficients, and $v_{i}^{(n)},u_{j}^{(n)}$ are row- and column-basis vectors. To promote a low-rank and spatially consistent solution, we introduce entropy and variance penalties.
@@
encouraging dominance of only a few modes to enforce low rank.
Additionally, to discourage memorization of individual indices, we penalize variance in $v^{(n)}$ and $u^{(n)}$:
@@
Applying these regularizations to the embedding matrices of both $x$- and $y$-factors results in more structured and factorized representations, improving generalization to unseen coordinate pairs. Fig. 3(b) illustrates that, without regularization, the model struggles to preserve signal integrity in OOD regions. With regularization, the embeddings exhibit structured banding aligned with coordinate axes, indicating a discovered factorized representation beneficial for compositional extrapolation. In addition, Fig. 3(c) show that with regularization, the learned representation is significantly smoother than without, despite having the same factorized embedding architecture. Both results show
@@
To assess the impact of our regularization strategy, we compare four configurations: (1) No Reg (baseline MSE loss), (2) Entropy Only (applying Eq. (10)), (3) Variance Only (applying Eq. (11)), and (4) Entropy+Var (combining both). Fig. 3(d) plots in-distribution (ID) vs. out-of-distribution (OOD) MSE. The No Reg condition yields the lowest ID error but highest OOD error, reflecting poor generalization. Entropy Only marginally improves OOD performance, while Variance Only significantly enhances extrapolation at the cost of higher ID loss. Combining both constraints (Entropy+Var) achieves the best OOD performance while maintaining low ID error, striking the optimal balance for robust factorization.
Intuitively, the regularization leverages simple, low-dimensional structure in the learned embedding matrices. By limiting the effective rank, it forces the network to represent each factor (e.g. a coordinate dimension) through a small set of shared directions rather than separately memorizing every possible input combination. Consequently, once the model learns an appropriate low-rank basis for in-distribution data, it can more easily compose those basis elements to handle new, unseen factor combinations in the OOD region. In other words, restricting the embedding to a few dominant modes discourages the network from relying on specific entries for every $(x,y)$, thereby reducing overfitting and promoting a genuinely factorized representation that better generalizes to novel coordinates.
@@ 4.2 Dataset Augmentation
Another approach that successfully enabled compositional generalization in the toy setting is dataset curation. Providing independent factors of variation in the form of x “stripes” and separate y stripes help the network discover a representation that enables composition.
Concretely, we generate 1D vertical and horizontal stripes by fixing a single coordinate (either $x$ or $y$) and applying a 1D Gaussian in the orthogonal direction (samples shown in Fig. 4(a), full set shown in Fig. 7(a)). These stripes in the full output representational space act as building blocks that allow the network to learn each factor of variation independently in a way that abstract factorized inputs were not able to do. Optionally, we additionally include a small number of the 2D bumps (outside the OOD region, as before) in the training set, so the model also sees a small number of 2D examples. Consequently, the network acquires strong compositional ability: it generalizes to novel $(x,y)$-combinations outside the original training distribution (even with zero 2D bumps). Empirically, as demonstrated in Fig. 4, this approach markedly improves extrapolation performance.
@@
Remarkably, we observed that models trained on a dataset consisting of just the stripes are able to compositionally generalize to generate additive stripe conjunctions in the OOD region without having seen any compositional examples (inside or outside the OOD region) during training, as shown in Fig. 4(b) when 0 2D Gaussian bump images are included in the training dataset. The remaining panels of Fig. 4(b) show that as we increase the number of 2D Gaussian bump images included in the training dataset, the model learns to generate proper bumps in the OOD region (also see Fig. 7(b)). We further characterize the data efficiency of such an approach of learning to compositionally generate bumps in the OOD region (learning with stripes + bumps) versus a traditional approach of learning with just bumps. Here we benchmark the compared models based on the minimal threshold amount of data needed to reach a target accuracy of $90\%$ in generating the correct $x$ and $y$ coordinates of the stripe conjunctions/bumps. To our surprise, we found that models trained on datasets consisting of purely bumps has a data threshold that scales as the cube of the per-side image size, hence data requirements grow as the cube of the problem size, $N$. On the other hand, due to the zero-shot generalization capability of models trained on the stripe dataset, the scaling is linear with $N$, as the dataset consists of $2N$ stripes in total, $N$ for $x$ and $N$ for $y$ (together with 0 or a small number of bumps).
@@
For comparison with the earlier results, we again plotted the volume metrics for the model trained on stripes + bumps dataset, as a function of model layer depth. The representations remain relatively smooth and uniform even across the OOD region, in contrast to the models trained with just the bumps shown in Fig. 2. In addition, we show in Fig. 4(e) the 2D activations of neurons in the first layer of a network trained on a dataset of stripes + 50 bumps. The activations show consistent “stripe” patterns that resemble the embeddings shown in Fig. 3(b) with regularization. Moreover, we show neural tuning curves w.r.t. to $x$ and $y$ for two sample neurons in layer 0 and layer 3 of the same network in Fig. 4(f), which exhibits “stripes” in early layers, which localizes to become a “bump” in downstream layers (more sample neural tuning curve can be found in Fig. 15).
@@
Intuitively, these stripes provide a “scaffold” in the output space that is straightforward for the network to learn via additive composition. The model then leverages this scaffold to handle the more complex multiplicative compositional task—namely, generating intersections instead of mere unions—thereby achieving efficient bump generalization in the compositional OOD region. This highlights the value of data-centric strategies that emphasize the generative factors in isolation, consistent with the finding that dataset augmentation with stripes also improves compositional generalization and data-efficiency scaling for 2D bump generation in diffusion models in Ref. (Liang et al., 2024b).
@@
The commonality in the two disparate methods that enable compositional generalization is that they both encourage factorization directly at the level of the 2D image space. By contrast, even directly supplying the network with factorized representations does not work if those representations are abstract and in a different space than the 2D image space. Our findings indicate that factorization learned in a latent space may not survive through subsequent dimension expansion in a decoder. Instead, to achieve robust compositional generalization, we found that it was important to explicitly build or demonstrate factorization represented in the model’s output space. By ensuring that the network learns to separate each underlying factor (e.g. $x$ and $y$) at the pixel level, we preserve the necessary compositional structure for strong extrapolation to novel conditions.
@@ 5 Discussion & Conclusion
    Disentangled latents are not enough. Simply factorizing a single layer (e.g. the bottleneck) does not guarantee compositional structure in deeper layers; manifold warping can re-entangle factors and fail in OOD regions.
@@
    Enforcing factorization in the output space is key. Through explicit low-rank regularization or domain-centric augmentation (adding 1D stripes), we can maintain separable $x$/$y$ embeddings and achieve strong extrapolation. In other words, building or preserving factorized structure across the entire model forward pass is essential for truly compositional generalization.
@@
While our investigation provides a clean, mechanistic view of why disentangled representations often fail on a synthetic 2D bump task, its scope is inherently limited by the toy nature of the setting. Extending these insights to large-scale models (e.g. Transformers, auto-regressive architectures) and high-dimensional data (vision, language) is an important next step to see whether similar re-entanglement dynamics persist. Furthermore, modern architectures like vision transformers or large language models may already embed implicit factorization biases (e.g. through attention or prompt-based modularity); rigorously assessing whether these biases genuinely promote compositional generalization would sharpen our understanding of inductive bias design. Finally, new metrics or frameworks are needed to capture partial and hierarchical compositional structures, since real-world tasks rarely exhibit a single, strict factorization. We provide a preliminary attempt for characterizing such partial factorization in Appendix A. By developing principled ways to quantify such subtler forms of compositionality, future work can more precisely guide architectural choices and regularization strategies toward robust OOD performance in more complex, real-world domains.
@@ B.1 General $N$-Dimensional Definition
$\mathrm{d}v(\mathbf{x})\;=\;\sqrt{\det\bigl{(}g(\mathbf{x})\bigr{)}}\;\mathrm{d}x_{1}\,\dots\,\mathrm{d}x_{N}.$ (39)
@@ Appendix C Low-rank Regularization for Compositional Generalization
$\mathcal{L}_{\text{ent}}\;=\;-\,\eta_{1}\,\sum_{n=1}^{r}\widetilde{\lambda}^{(n)}\,\ln\bigl{[}\widetilde{\lambda}^{(n)}\bigr{]}.$ (46)
@@
Appendix D Dataset Augmentation

### v2

@@ 4 Encouraging Compositional Generalization
Figure 3: Inducing compositional generalization through architectural rendering and regularization constraints. (a) Schematic of architecturally forced rendering of the initially disentangled representations (with 1-hot input encodings) into a space matching the output space (disentangled processing). (b) Sampled embedding activations corresponding to $x=14$ and $y=14$ for networks trained without and with regularization, respectively. (c) Generated images when the OOD region is on the top right corner, for the non-regularised (top) and regularised networks (bottom) respectively. (d) Volume metric comparison between networks trained without and with regularization as a function of layer depth, respectively. (e) OOD vs. ID MSE plot for various ablation studies over many runs.
Figure 4: Inducing generalizable composition by training single disentangled factors to render (data curriculum). (a) Sample grayscale images of 1D Gaussian “stripes” at $x=14$ and $y=14$. (b) Generated OOD output as a function of number of Gaussian bumps (all outside the upper right OOD region) included in the training dataset. (c) Data/sample efficiency: data scaling of the stripes-only ($\sim N$) vs. bumps-only ($\sim N^{3}$) datasets to reach $90\%$ accuracy of $x$ and $y$ generation as a function of image size $N$ (accuracy assessed based on the location of the darkest pixel). Here the stripes + bumps dataset consists of $\sim 7N$ bumps + $2N$ stripes; learning breaks the curse of dimensionality due to the ability to compositionally generalize zero-shot. (d) Volume metric as a function of layer depth of the network trained on a dataset of stripes + bumps. (e) 2D neuron activations across different channels in the first layer (layer 0) of a network trained trained on a dataset of stripes + 50 bumps. (f) Neural tuning curves of two sample neurons at layer 0 and layer 3 of the same network as in (e).
@@
4.1 Architecture regularization for low-rank embeddings
One approach we adopt is to architecturally encourage the expansion of abstract latent representations into structured 2D embedding filters, while applying explicit rank regularization during training (Fig. 3(a)). This design promotes interpretability and compositional disentanglement in the learned features. For example, Cichocki et al. (2009) demonstrate that enforcing low-rank tensor factorization can mitigate spurious interactions between factors. Inspired by this, we apply a matrix (or tensor) factorization penalty to the set of 2D embedding filters associated with each input dimension, encouraging them to remain approximately low-rank and spatially consistent.
To formalize this, in the 1-hot input encoding setup, we generate embedding matrices of shape $N\crossproduct N$ for each token $x$ and separately for $y$ (Fig. 3(a)). There are a total of $2N$ such matrices. Denoting one of them as $W_{i,j}$, we decompose this matrix using an SVD-like form:
@@ 4.1 Architecture regularization for low-rank embeddings
where $r$ is a chosen upper bound on the rank of the decomposition, $\lambda^{(n)}$ are scalar singular values, and $v_{i}^{(n)}$, $u_{j}^{(n)}$ are orthogonal basis vectors along the row and column dimensions, respectively. This formulation enables us to directly regularize the singular value spectrum and the spatial smoothness of basis vectors, promoting low-rank structure in each embedding matrix.
This regularization is particularly effective because each input dimension is associated with a dedicated 2D embedding matrix, which constrains the model to learn spatially structured factors aligned with individual input dimensions. In practice, we observe that this encourages the model to discover localized or “stripe-like” patterns in the learned embeddings (see Fig. 3(b)), facilitating compositional generalization. To further stabilize learning and reduce redundancy, we additionally penalize the entropy and variance of the singular values, which promotes sparsity and consistent usage of the factorized components.
@@
encouraging dominance of only a few modes to enforce low rank. Additionally, to discourage memorization of individual indices, we penalize variance in $v^{(n)}$ and $u^{(n)}$:
@@
Applying these regularizations to the embedding matrices of both $x$- and $y$-factors results in more structured and factorized representations, improving generalization to unseen coordinate pairs. Fig. 3(b) illustrates that, without regularization, the model struggles to preserve signal integrity in OOD regions. With regularization, the embeddings exhibit structured banding aligned with coordinate axes, indicating a discovered factorized representation beneficial for compositional extrapolation. In addition, Fig. 3(c) shows that our regularization works even for a corner OOD region, indicating a stronger form of (extrapolative) compositional generalization. With regularization, the learned representation is significantly smoother than without, as evidenced in Fig. 3(d), despite having the same factorized embedding architecture. These results demonstrate the effectiveness of our proposed regularization technique.
@@
To assess the impact of our regularization strategy, we compare four configurations: (1) No Reg (baseline MSE loss), (2) Entropy Only (applying Eq. (10)), (3) Variance Only (applying Eq. (11)), and (4) Entropy+Var (combining both). Fig. 3 (e) plots MSE in distribution (ID) versus out-of-distribution (OOD). The condition No Reg yields the lowest ID error but the highest OOD error, reflecting poor generalization. Entropy Only marginally improves OOD performance, while Variance Only significantly enhances extrapolation at the cost of higher ID loss. Combining both constraints (Entropy+Var) achieves the best OOD performance while maintaining low ID error, striking the optimal balance for robust factorization.
Intuitively, the regularization leverages simple, low-dimensional structure in the learned embedding matrices. By limiting the effective rank, it forces the network to represent each factor (for example, a coordinate dimension) through a small set of shared directions rather than separately memorizing every possible input combination. Consequently, once the model learns an appropriate low-rank basis for in-distribution data, it can more easily compose those basis elements to handle new, unseen factor combinations in the OOD region. In other words, restricting the embedding to a few dominant modes discourages the network from relying on specific entries for every $(x,y)$, thus reducing overfitting and promoting a genuinely factorized representation that better generalizes to novel coordinates.
@@ 4.2 Dataset augmentation
Another approach that successfully enabled compositional generalization in the toy setting is dataset curation. Asking the network to generate independent factors of variation in the form of separate $x$ and $y$ “stripes” help the network discover a representation that enables compositionality.
Concretely, we generate 1D vertical and horizontal stripes by fixing a single coordinate (either $x$ or $y$) and applying a 1D Gaussian in the orthogonal direction (samples shown in Fig. 4(a), full set shown in Fig. 7(a)). Note that here the 2D input structure to the model is still maintained via setting the coordinate to the fixed dimension to $-1$. These stripes in the full output representational space act as building blocks that allow the network to learn each factor of variation independently in a way that abstract factorized inputs are unable to. Optionally, we additionally include a small number of the 2D bumps (outside the OOD region, as before) in the training set, so the model also sees a small number of 2D examples. Consequently, the network acquires strong compositional ability: it generalizes to novel $(x,y)$-combinations outside the original training distribution (even with zero 2D bumps). Empirically, as demonstrated in Fig. 4, this approach markedly improves compositional generalization performance.
@@
Remarkably, we observed that models trained on a dataset consisting of just the stripes are able to compositionally generalize to generate additive stripe conjunctions in the OOD region without having seen any compositional examples (inside or outside the OOD region) during training, as shown in Fig. 4(b) when 0 2D Gaussian bump images are included in the training dataset. Similar to Fig. 3(c), we show here an example of corner OOD region to demonstrate that our method is robust for extrapolative compositional generalization. The remaining panels of Fig. 4(b) show that as we increase the number of 2D Gaussian bump images included in the training dataset, the model learns to generate proper bumps in the OOD region (also see Fig. 7(b)). We further characterize the data efficiency of this approach to learning to compositionally generate bumps in the OOD region (learning with stripes + bumps) versus a traditional approach to learning with just bumps. The total data count scaling with respect to the image size $N$ is shown in Fig. 4(c). Here we benchmark the compared models based on the minimal threshold amount of data needed to reach a target accuracy of $90\%$ in generating the correct $x$ and $y$ coordinates of the stripe conjunctions/bumps. To our surprise, we found that models trained on datasets consisting of purely bumps have a data threshold that scales as $\sim N^{3}$, and hence the data requirements grow as the cube of the size of the problem. On the other hand, due to the zero-shot generalization capability of models trained on the stripe dataset, the scaling is linear with $N$, since the data set consists of $2N$ stripes in total, $N$ for $x$ and $N$ for $y$ (together with 0 or a small number of bumps). Furthermore, with the increase of stripes training with bumps to obtain bump outputs rather than plus-shaped ones, the minimum required number of training bumps scales only linearly with $N$, meaning that the total training data remains linear in $N$, indicative of compositional generalization.
@@
For comparison with the earlier results, we again plotted the volume metrics for the model trained on the stripes + bumps dataset, as a function of model layer depth, Fig. 4(d). The representations remain relatively smooth and uniform even in the OOD region, contrasting the models trained with just the bumps shown in Fig. 2. In addition, we show in Fig. 4(e) the 2D activations of neurons in the first layer of a network trained on a dataset of stripes + 50 bumps. The activations show consistent “stripe” patterns that resemble the embeddings shown in Fig. 3(b) with regularization. Moreover, we show neural tuning curves w.r.t. to $x$ and $y$ for two sample neurons in layer 0 and layer 3 of the same network in Fig. 4(f), which exhibits “stripes” in early layers, which localizes to become a “bump” in downstream layers (more sample neural tuning curve can be found in Fig. 15).
@@
Intuitively, the early-layer stripes allow the network to build a “scaffold” in its representations that are straightforward for it to use to perform additive composition. Later layers in the model then leverage this scaffold to handle the multiplicative compositional task — to generate stripe intersections instead of unions — thereby achieving efficient bump generalization in the OOD region. This highlights the value of data-centric strategies that emphasize the generative factors in isolation, consistent with the finding that dataset augmentation with stripes also improves compositional generalization and data-efficiency scaling for 2D bump generation in diffusion models in Ref. (Liang et al., 2024b). While the prior work concluded that stripe training generated disentangled latent representations, the present work shows that supplying or learning disentangled latents is not sufficient for compositional generalization, and that stripe training even with disentangled latents helps form pixel-space latents representations that are necessary for OOD compositional generalization.
@@
The commonality between the two disparate approaches that enable OOD generalization is that they both encourage the network to form disentangled representations that are encoded in the space of the 2D pixel outputs. By contrast, even directly supplying the network with factorized representations does not work if those representations are abstract and in a different latent space than the 2D image space. Factorization in a latent space may not survive subsequent dimension expansion by a downstream decoder. By ensuring that the network learns to separate each underlying factor (e.g. $x$ and $y$) at the pixel level, we preserve the necessary compositional structure for strong extrapolation to novel conditions.
@@ 5 Discussion & Conclusion
    Disentangled latents are not enough. Factorizing a single latent layer (e.g. the bottleneck) into a disentangled representation does not guarantee compositional structure in deeper layers; manifold warping can re-entangle factors, leading to failure in OOD regions.
@@
    Rendering disentangled latents into the same representational space as the outputs is key. Through a specific architecture with low-rank regularization or domain-centric data curation (training with independent factors), the network generates pixel-level embeddings of the individual $x$, $y$ latents to achieve strong compositional generalization.
@@
While our investigation provides a clean, mechanistic insight into why disentangled representations often fail on a synthetic 2D bump task, its scope is inherently limited by the toy nature of the setting. Extending these insights to large-scale models (e.g., diffusion models, autoregressive architectures, Transformers) and high-dimensional data domains (e.g., vision, language) is an important next step. Modern architectures like vision transformers or large language models may already embed implicit factorization biases—through attention mechanisms or prompt-based modularity—but rigorously assessing whether these biases genuinely support compositional generalization remains an open challenge.
We also acknowledge that translating insights from our synthetic study to more complex, real-world datasets is nontrivial but essential. As highlighted by Montero et al. (2022), interactive compositionality is unlikely to be addressed by a one-size-fits-all solution. Accordingly, we do not claim our approach is universally applicable, but rather that it reveals two promising and generalizable directions: (1) training modular, output-level embedding filters dedicated to each disentangled input dimension, and (2) leveraging dataset augmentation strategies that isolate factors of variation. Our simplified setting was deliberately chosen to cleanly expose the underlying failure modes, such as the model memorizing ID configurations for OOD generalization, without confounding interactions.
In future work, we plan to explore how these principles scale to standard disentanglement datasets and more naturalistic settings, explicitly investigating whether modular embedding and low-rank factorization constraints retain their utility. Finally, we emphasize the need for new metrics and frameworks to characterize partial and hierarchical compositional structures, which are far more representative of real-world cognition. We provide a preliminary attempt at capturing such structures in Appendix A. Developing principled tools to measure subtler forms of compositionality may better guide both architectural and algorithmic choices in the pursuit of robust generalization in complex domains.
@@ Acknowledgements
Impact Statement
Our work develops theoretical and empirical insights into why disentangled or factorized inputs do not necessarily yield compositional generalization, and proposes practical strategies (data augmentation, low-rank regularization) to preserve factorization within the network. By illuminating how and why factorization can break down—or be maintained—this research can guide future model designs that are more transparent, data-efficient, and capable of robust extrapolation beyond the training distribution.
From an ethical or societal standpoint, we do not foresee immediate risks arising directly from these factorization techniques: the methods and analyses we present are primarily at the conceptual and architectural level. However, we note that improvements in compositional generalization can enable models to be more adaptable, including in sensitive real-world domains such as medical imaging or autonomous systems. While better extrapolation is desirable, it also entails that models could behave in unintended ways under distribution shifts if they mix factors in novel, untested combinations. Ultimately, we believe our framework, by clarifying the conditions needed for truly factorized representations, will enhance the interpretability and reliability of neural networks, a net positive for the broader machine learning community.
@@ B.1 General $N$-Dimensional Definition
$\mathrm{d}v(\mathbf{x})\;=\;\sqrt{\det\bigl(g(\mathbf{x})\bigr{missing})}\;\mathrm{d}x_{1}\,\dots\,\mathrm{d}x_{N}.$ (39)
@@ Appendix C Low-rank Regularization for Compositional Generalization
$\mathcal{L}_{\text{ent}}\;=\;-\,\eta_{1}\,\sum_{n=1}^{r}\widetilde{\lambda}^{(n)}\,\ln[\widetilde{\lambda}^{(n)}\bigr{]}.$ (46)
@@
Appendix D Experimental Details
In this section, we describe the neural network architectures used for generating bump functions, including exact layer sizes, activation functions, optimization algorithms, hyperparameters, and training protocols. Code to reproduce the experiments can be found in this Github repository. We explored various input formats and model depths; below are detailed descriptions of representative CNN and MLP models using bump-encoding inputs:
Property CNN Architecture MLP Architecture
Input Dimensions 56 (reshaped to $56\times 1\times 1$) 56
Output Dimensions $1\times 28\times 28$ $1\times 28\times 28$
Layers 4 ConvTranspose2d layers (upsampling from $7\times 7$ to $28\times 28$) 3 fully-connected linear layers
Hidden Layer Size 64 channels per hidden layer 256 units per hidden layer
Parameter Count $\sim$315K $\sim$272K
Activation Function ReLU ReLU
Optimizer AdamW (Learning Rate: $1\times 10^{-3}$) AdamW (Learning Rate: $1\times 10^{-3}$)
Table 1: Comparison of CNN and MLP architectures
These details clearly communicate the architectural specifics and ensure reproducibility.
Appendix E Dataset Augmentation